# Unified Analysis of Screening Masses for Vector and Axial-Vector Mesons and Their Diquark Partners in the Contact Interaction Model

### v1

@@ I Introduction
From the theoretical perspective, while the thermal evolution of hadronic states in the light-quark sector has been extensively investigated, see Refs. Chen et al. (2024); Mukherjee (2009); Dosch and Narison (1988); Ayala et al. (2012); Hatsuda et al. (1993); Wang and Wang (2015), studies of the heavy-quark sector remain relatively scarce, despite their critical importance. Quarkonia formation and dissociation provide clear signatures of the QGP, making the precise determination of their screening masses essential for heavy-ion collision phenomenology Brambilla and others (2011). Furthermore, quarkonium suppression has long been proposed as an indicator of deconfinement Matsui and Satz (1986). Thus, exploring the behavior of quarkonium under extreme conditions is directly tied to understanding the properties of strongly interacting matter in the early universe.
Mesons, as bound states of two quarks—either light or heavy—play a central role in these investigations. In fact, it is well established that the first phenomenological order parameter for quark–gluon deconfinement was introduced in the vector (V) channel ($J^{PC}=1^{--}$) through thermal QCD sum rules, as reported in Ref. Bochkarev and Shaposhnikov (1986). This pioneering work was followed by numerous studies Dosch and Narison (1988); Ayala et al. (2012); Hatsuda et al. (1993); Wang and Wang (2015), and a broad spectrum of theoretical models has since been developed to describe V mesons at finite temperature. These approaches include effective chiral Lagrangians Song (1996), the sigma model Pisarski (1995), the Kroll–Lee–Zumino model Hernández et al. (2025), the Schwinger-Dyson and Bethe–Salpeter equations (BSE) framework Gao and Ding (2020); Fischer (2019), the nonlocal Polyakov–Nambu–Jona-Lasinio model Carlomagno and Izzo Villafañe (2019), Gribov quantization Sumit et al. (2023), and lattice QCD simulations Cheng and others (2011); Mukherjee (2009); Bazavov and Weber (2021); Petreczky (2009), among others.
The axial-vector (AV) channel has also received considerable attention. The lightest meson in this sector, the $a_{1}$, has been analyzed through thermal QCD sum rules to study its evolution at finite temperature Dominguez et al. (2013b); Hatsuda et al. (1993); Dey et al. (1990); Mamedov and Taghiyeva (2021). In addition, the thermal behavior of heavy V states has been examined: the $J/\Psi$ in Ref.Dominguez et al. (2010), ground-state bottomonia ($\Upsilon$) in Ref.Dominguez et al. (2013a), and V charmonia and bottomonia through bottom-up AdS/QCD approaches in Ref. Martin Contreras et al. (2021).
@@
A natural extension of meson studies is the investigation of diquark correlations, which are central to the description of multiquark bound states, particularly baryons in the quark–diquark picture Gell-Mann (1962); Ida and Kobayashi (1966); Lichtenberg and Tassie (1967). Diquarks are color–non-singlet correlations of two quarks and, owing to their color charge, remain confined within hadrons such as baryons, tetraquarks, or pentaquarks, precluding direct experimental observation. Although diquarks have been widely studied at zero temperature in the context of hadron spectroscopy and nucleon structure Cahill et al. (1987, 1989); Oettel et al. (1998); Gutiérrez-Guerrero et al. (2019, 2021); Chen et al. (2012); Barabanov and others (2021), their thermal properties are comparatively less understood. The evolution of diquark correlations with temperature provides insight into the possible persistence or dissolution of baryonic and tetraquark states in the medium, making this analysis a genuine complement to meson studies. Since diquark dynamics directly influence baryon survival and the behavior of hot QCD matter, their role at finite temperature is far from trivial Chen et al. (2024); Wang et al. (2013). In this work, which also considers diquarks containing heavy quarks, we aim to develop a consistent framework for determining the masses of light and heavy baryons—and possibly tetraquarks Bedolla et al. (2020)—at finite temperature. Our analysis lays the groundwork for future studies of hadrons under thermal conditions, where the relative contribution of each diquark channel, and its temperature dependence, is expected to play a critical role. Then, this as an initial step toward a systematic and unified understanding of hadronic matter at finite temperature.
@@ II.3 Quark Masses
We illustrate the evolution of the screening masses of these quarks as a function of temperature in Fig. 3. In the present work, which includes mesons containing heavy quarks, we adopt $T_{c}=155$ MeV, consistent with values used in lattice QCD simulations Andronic et al. (2018); Bazavov and others (2012), to facilitate comparison with other approaches. However, it is important to emphasize that this temperature acts merely as a rescaling factor in our approach.
FIG. 3: Results for the screening quark masses computed across a temperature range from 0 to $4\,T_{c}$, with $T_{c}=155$ MeV Andronic et al. (2018); Bazavov and others (2012). At high temperatures, the quark mass asymptotically tends toward its bare mass, $M(T)\to m$.
@@ III.1 Vector Mesons
In our analysis we consider spin-1 mesons, starting with V states built from heavy ($Q\bar{Q}$), heavy–light ($Q\bar{q}$), and light ($q\bar{q}$) configurations, focusing on the $\color[rgb]{1,0,0}u$}, $\color[rgb]{0.64,0.0,0.0}d$}, $\color[rgb]{0,0.5,0}s$}, $\color[rgb]{0.55,0.0,0.55}c$}, and $\color[rgb]{0,0,1}b$} flavors within the $SU(5)$ multiplets of the quark model Gell-Mann (1964); Zweig (1964b, a).
@@
It is evident that at zero temperature, our results can be compared with experimental data Navas and others (2024), and they show good agreement. At zero temperature, $E_{{\color[rgb]{0,0,1}V}}^{L}=E_{{\color[rgb]{0,0,1}V}}^{T}$ and $m_{{\color[rgb]{0,0,1}V}}^{L}=m_{{\color[rgb]{0,0,1}V}}^{T}$, so in this case the longitudinal and transverse labels are omitted. For our calculations involving mesons composed of different quark flavors, we modify Eq. (10) by replacing $M_{f}$ with the reduced mass $M_{R}=M_{{\color[rgb]{0,0,1}f_{1}}}M_{{\color[rgb]{0,0,1}\bar{f_{2}}}}/[M_{{\color[rgb]{0,0,1}f_{1}}}+M_{{\color[rgb]{0,0,1}\bar{f_{2}}}}]$. In the last column of the table we report the screening-mass limit at 500 MeV. In this limit, the masses of the light mesons display a percentage increase exceeding 100% with respect to their $T=0$ MeV values, whereas for the heavy mesons the increase is much smaller, about 6%.
FIG. 5: Longitudinal and transverse screening masses for V mesons composed of the five quark flavors($\color[rgb]{1,0,0}u$}, $\color[rgb]{0.64,0.0,0.0}d$}, $\color[rgb]{0,0.5,0}s$}, $\color[rgb]{0.55,0.0,0.55}c$} and $\color[rgb]{0,0,1}b$}). The longitudinal components are shown with dashed lines, while the transverse components are shown with solid lines. In the top left panel, the lightest vector meson is shown, the $\rho$ meson, The black dashed line corresponds to the free-theory limit, $m=2\pi T/T_{c}$; the top center panel displays the $K_{1}$ screening mass results; the top right panel displays strangeonium vector mesons. The bottom left panel is dedicated to vector $D$-mesons ($D^{*0}$ and $D^{*}_{s}$); the bottom center panel is committed for vector $B$-mesons ($B^{+*}$, $B^{0*}_{s}$ and $B^{*}_{c}$) and, the bottom right panel shows the results for charmonium and bottomonium vector mesons. Results shown with squares and diamonds correspond to lattice results from Refs. Bazavov and Weber (2021); Aarts et al. (2022), while stars and inverted triangles represent results obtained using an effective chiral Lagrangian from Refs. Montana et al. (2022); Song (1993). Pentagons and right triangles correspond to predictions obtained using QCD sum rules in Refs. Veliev et al. (2011); Dominguez et al. (2013a).
@@
In addition, we present in Fig. 5 a comparison of our results for the transverse component with lattice data from Refs. Bazavov and Weber (2021); Aarts et al. (2022), as well as with predictions obtained from an effective chiral Lagrangian model Montana et al. (2022); Song (1993) and from QCD sum rules Veliev et al. (2011); Dominguez et al. (2013a). In most models the screening mass increases with temperature, but in mesons containing one or two bottom quarks it may decrease, as predicted by QCD sum rules Ref. Veliev et al. (2011); Dominguez et al. (2013a) and effective chiral Lagrangian model Song (1993). The behavior of the screening masses in these two approaches is similar to our CI results at low temperatures; however, clear deviations appear once the system reaches and exceeds the critical temperature. This mainly originates from a limitation of finite-temperature QCD sum rules: correlators are evaluated only at leading order in perturbative QCD (one loop), and higher-order thermal corrections are not reliable. Therefore, these calculations are valid only for $0\leq T\leq T_{c}$, directly influencing the temperature dependence of the screening masses, as discussed in Ref. Dominguez et al. (2013a) and confirmed in Ref. Veliev et al. (2011), where it is explicitly stated that the results should not be extended to $T>T_{c}$. A similar decreasing trend is observed in the points marked with stars Song (1993). In that work, the results are only reliable near $T_{c}$, and for higher temperatures the predictions become ambiguous, since the model cannot describe the symmetry-restored regime. This variation among theoretical predictions highlights the fragmented state of the existing literature, particularly in the heavy-meson sector. Since the available results are derived from fundamentally different methodologies and assumptions, their temperature dependences cannot be compared on equal footing. Although we have compiled the broadest set of screening-mass results accessible, they necessarily originate from diverse approximations, which accounts for the discrepancies observed. The present work seeks to address this issue by offering a unified analysis within a single, consistent theoretical framework.
@@
Also in several lattice QCD studies, only the transverse component of the screening mass is reported. This choice is motivated by the fact that, at high temperature or nonzero chemical potential, the longitudinal propagator is strongly suppressed in the infrared, whereas the transverse one remains comparatively stable Bazavov and others (2019). Moreover, the transverse channel is usually determined with better numerical precision in lattice simulations, making it the most reliable observable. Thus, the results presented in this work provide the CI model prediction for the transverse component over a broader temperature range than for the longitudinal one, which is calculated only up to $T_{c}$. This extension constitutes a valuable step toward a more complete description, as it allows for analytical expressions of screening masses and BSAs. However, the same analytical treatment also increases the complexity of incorporating longitudinal effects beyond $T_{c}$, making predictions above this value less robust. Nevertheless, the present results offer a consistent baseline that can guide future refinements and extensions of the approach.
@@ III.2 Axial-Vector Mesons
FIG. 6: Longitudinal and transverse screening masses for AV mesons composed of the five quark flavors($\color[rgb]{1,0,0}u$}, $\color[rgb]{0.64,0.0,0.0}d$}, $\color[rgb]{0,0.5,0}s$}, $\color[rgb]{0.55,0.0,0.55}c$} and $\color[rgb]{0,0,1}b$}).The longitudinal components are shown with dashed lines, while the transverse components are shown with solid lines. In the top left panel, the lightest axial-vector meson is shown, the $a_{1}$ meson, The black dashed line corresponds to the free-theory limit, $m=2\pi T/T_{c}$; the top center panel displays the $K_{1}^{*}$ screening mass results; the top right panel displays strangeonium axial-vector mesons. The bottom left panel is dedicated to axial-vector $D$-mesons ($D_{1}$ and $D_{{\color[rgb]{0,0.5,0}s}1}$); the bottom center panel is committed for axial-vector $B$-mesons ($B_{1}$, $B_{{\color[rgb]{0,0.5,0}s}1}$ and $B^{*}_{{\color[rgb]{0.55,0.0,0.55}c}{\color[rgb]{0,0,1}b}}$) and, the bottom right panel shows the results for charmonium and bottomonium axial-vector mesons. Squares and crosses denote lattice results from Refs. Bazavov and Weber (2021); Cheng and others (2011), inverted triangles represent results from an effective chiral Lagrangian in Ref. Song (1993), and hexagons indicate QCD sum-rule predictions from Ref. Yazici (2016).
@@
Figure 6 illustrates the temperature dependence of the screening masses over the range $T/T_{c}\in[0,250]$ MeV. Transverse results are shown with solid lines, longitudinal results with dashed lines. Analogously to the vector mesons, we show the free-theory limit with a black dashed line in the case of light mesons. Our results are compared with lattice data Bazavov and Weber (2021); Cheng and others (2011); Mukherjee (2009), an effective chiral Lagrangian Song (1993), and QCD sum-rule predictions Yazici (2016). It should be stressed that the results derived from the last two approaches are reliable only for temperatures lower than the critical temperature, for the same reasons explained for V mesons.

### v2

@@ I Introduction
From the theoretical perspective, while the thermal evolution of hadronic states in the light-quark sector has been extensively investigated, see Refs. Chen et al. (2024); Mukherjee (2009); Dosch and Narison (1988); Ayala et al. (2012); Hatsuda et al. (1993); Wang and Wang (2015), studies of the heavy-quark sector remain relatively scarce, despite their critical importance. Quarkonia formation and dissociation provide clear signatures of the QGP, making the precise determination of their screening masses essential for heavy-ion collision phenomenology Brambilla et al. (2011). Furthermore, quarkonium suppression has long been proposed as an indicator of deconfinement Matsui and Satz (1986). Thus, exploring the behavior of quarkonium under extreme conditions is directly tied to understanding the properties of strongly interacting matter in the early universe.
Mesons, as bound states of two quarks—either light or heavy—play a central role in these investigations. In fact, it is well established that the first phenomenological order parameter for quark–gluon deconfinement was introduced in the vector (V) channel ($J^{PC}=1^{--}$) through thermal QCD sum rules, as reported in Ref. Bochkarev and Shaposhnikov (1986). This pioneering work was followed by numerous studies Dosch and Narison (1988); Ayala et al. (2012); Hatsuda et al. (1993); Wang and Wang (2015), and a broad spectrum of theoretical models has since been developed to describe V mesons at finite temperature. These approaches include effective chiral Lagrangians Song (1996), the sigma model Pisarski (1995), the Kroll–Lee–Zumino model Hernández et al. (2025), the Schwinger-Dyson and Bethe–Salpeter equations (BSE) framework Gao and Ding (2020); Fischer (2019), the nonlocal Polyakov–Nambu–Jona-Lasinio model Carlomagno and Izzo Villafañe (2019), Gribov quantization Sumit et al. (2023), and lattice QCD simulations Cheng et al. (2011); Mukherjee (2009); Bazavov and Weber (2021); Petreczky (2009), among others.
The axial-vector (AV) channel has also received considerable attention. The lightest meson in this sector, the $a_{1}$, has been analyzed through thermal QCD sum rules to study its evolution at finite temperature Dominguez et al. (2013a); Hatsuda et al. (1993); Dey et al. (1990); Mamedov and Taghiyeva (2021). In addition, the thermal behavior of heavy V states has been examined: the $J/\Psi$ in Ref.Dominguez et al. (2010), ground-state bottomonia ($\Upsilon$) in Ref.Dominguez et al. (2013b), and V charmonia and bottomonia through bottom-up Anti-de Sitter(AdS)/QCD approaches in Ref. Martin Contreras et al. (2021).
@@
A natural extension of meson studies is the investigation of diquark correlations, which are central to the description of multiquark bound states, particularly baryons in the quark–diquark picture Gell-Mann (1962); Ida and Kobayashi (1966); Lichtenberg and Tassie (1967). Diquarks are color–non-singlet correlations of two quarks and, owing to their color charge, remain confined within hadrons such as baryons, tetraquarks, or pentaquarks, precluding direct experimental observation. Although diquarks have been widely studied at zero temperature in the context of hadron spectroscopy and nucleon structure Cahill et al. (1987, 1989); Oettel et al. (1998); Gutiérrez-Guerrero et al. (2019, 2021); Chen et al. (2012); Barabanov et al. (2021), their thermal properties are comparatively less understood. The evolution of diquark correlations with temperature provides insight into the possible persistence or dissolution of baryonic and tetraquark states in the medium, making this analysis a genuine complement to meson studies. Since diquark dynamics directly influence baryon survival and the behavior of hot QCD matter, their role at finite temperature is far from trivial Chen et al. (2024); Wang et al. (2013). In this work, which also considers diquarks containing heavy quarks, we aim to develop a consistent framework for determining the masses of light and heavy baryons—and possibly tetraquarks Bedolla et al. (2020)—at finite temperature. Our analysis lays the groundwork for future studies of hadrons under thermal conditions, where the relative contribution of each diquark channel, and its temperature dependence, is expected to play a critical role. Then, this as an initial step toward a systematic and unified understanding of hadronic matter at finite temperature.
@@ II.3 Quark Masses
We illustrate the evolution of the screening masses of these quarks as a function of temperature in Fig. 3. In the present work, which includes mesons containing heavy quarks, we adopt $T_{c}=155$ MeV, consistent with values used in lattice QCD simulations Andronic et al. (2018); Bazavov et al. (2012), to facilitate comparison with other approaches. However, it is important to emphasize that this temperature acts merely as a rescaling factor in our approach.
FIG. 3: Results for the screening quark masses computed across a temperature range from 0 to $4\,T_{c}$, with $T_{c}=155$ MeV Andronic et al. (2018); Bazavov et al. (2012). At high temperatures, the quark mass asymptotically tends toward its bare mass, $M(T)\to m$.
@@ III.1 Vector Mesons
In our analysis we consider spin-1 mesons, starting with V states built from heavy ($Q\bar{Q}$), heavy–light ($Q\bar{q}$), and light ($q\bar{q}$) configurations, focusing on the $\color[rgb]{1,0,0}u$}, $\color[rgb]{0.64,0.0,0.0}d$}, $\color[rgb]{0,0.5,0}s$}, $\color[rgb]{0.55,0.0,0.55}c$}, and $\color[rgb]{0,0,1}b$} flavors within the $SU(5)$ multiplets of the quark model Gell-Mann (1964); Zweig (1964a, b).
@@
It is evident that at zero temperature, our results can be compared with experimental data Navas et al. (2024), and they show good agreement. At zero temperature, $E_{{\color[rgb]{0,0,1}V}}^{L}=E_{{\color[rgb]{0,0,1}V}}^{T}$ and $m_{{\color[rgb]{0,0,1}V}}^{L}=m_{{\color[rgb]{0,0,1}V}}^{T}$, so in this case the longitudinal and transverse labels are omitted. For our calculations involving mesons composed of different quark flavors, we modify Eq. (10) by replacing $M_{f}$ with the reduced mass $M_{R}=M_{{\color[rgb]{0,0,1}f_{1}}}M_{{\color[rgb]{0,0,1}\bar{f_{2}}}}/[M_{{\color[rgb]{0,0,1}f_{1}}}+M_{{\color[rgb]{0,0,1}\bar{f_{2}}}}]$. In the last column of the table we report the screening-mass limit at 500 MeV. In this limit, the masses of the light mesons display a percentage increase exceeding 100% with respect to their $T=0$ MeV values, whereas for the heavy mesons the increase is much smaller, about 6%.
FIG. 5: Longitudinal and transverse screening masses for V mesons composed of the five quark flavors($\color[rgb]{1,0,0}u$}, $\color[rgb]{0.64,0.0,0.0}d$}, $\color[rgb]{0,0.5,0}s$}, $\color[rgb]{0.55,0.0,0.55}c$} and $\color[rgb]{0,0,1}b$}). The longitudinal components are shown with dashed lines, while the transverse components are shown with solid lines. In the top left panel, the lightest vector meson is shown, the $\rho$ meson, The black dashed line corresponds to the free-theory limit, $m=2\pi T/T_{c}$; the top center panel displays the $K_{1}$ screening mass results; the top right panel displays strangeonium vector mesons. The bottom left panel is dedicated to vector $D$-mesons ($D^{*0}$ and $D^{*}_{s}$); the bottom center panel is committed for vector $B$-mesons ($B^{+*}$, $B^{0*}_{s}$ and $B^{*}_{c}$) and, the bottom right panel shows the results for charmonium and bottomonium vector mesons. Results shown with squares correspond to lattice-QCD screening-mass calculations from Ref. Bazavov and Weber (2021), while stars denote results obtained using an effective chiral Lagrangian approach at finite temperature from Ref. Montana et al. (2022). In channels where screening-mass results are not available, we compare our results with finite-temperature meson masses reported in the literature. In particular, lattice-QCD determinations and continuum-model calculations from Refs. Dominguez et al. (2013b); Aarts et al. (2022); Song (1993); Veliev et al. (2011) are shown using pentagons, diamonds, inverted triangles, and right-pointing triangles. These latter results do not correspond to screening masses but are included as qualitative benchmarks, providing complementary insight into the thermal behavior of vector mesons in sectors where direct screening-mass calculations are presently unavailable.
@@
In addition, we present in Fig. 5 a comparison of our results for the transverse component with lattice data from Refs. Bazavov and Weber (2021); Aarts et al. (2022), as well as with predictions obtained from an effective chiral Lagrangian model Montana et al. (2022); Song (1993) and from QCD sum rules Veliev et al. (2011); Dominguez et al. (2013b). In most models the mass increases with temperature, but in mesons containing one or two bottom quarks it may decrease, as predicted by QCD sum rules Ref. Veliev et al. (2011); Dominguez et al. (2013b) and effective chiral Lagrangian model Song (1993). The behavior of the masses in these approaches is similar to our CI results at low temperatures; however, clear deviations appear once the system reaches and exceeds the critical temperature. This mainly originates from a limitation of finite-temperature QCD sum rules: correlators are evaluated only at leading order in perturbative QCD (one loop), and higher-order thermal corrections are not reliable. Therefore, these calculations are valid only for $0\leq T\leq T_{c}$, directly influencing the temperature dependence of the screening masses, as discussed in Ref. Dominguez et al. (2013b) and confirmed in Ref. Veliev et al. (2011), where it is explicitly stated that the results should not be extended to $T>T_{c}$. A similar decreasing trend is observed in the points marked with stars Song (1993). In that work, the results are only reliable near $T_{c}$, and for higher temperatures the predictions become ambiguous, since the model cannot describe the symmetry-restored regime. This variation among theoretical predictions highlights the fragmented state of the existing literature, particularly in the heavy-meson sector. Since the available results are derived from fundamentally different methodologies and assumptions, their temperature dependences cannot be compared on equal footing. Although we have compiled the broadest set of screening-mass results accessible, they necessarily originate from diverse approximations, which accounts for the discrepancies observed. The present work seeks to address this issue by offering a unified analysis within a single, consistent theoretical framework.
@@
Also in several lattice QCD studies, only the transverse component of the screening mass is reported. This choice is motivated by the fact that, at high temperature or nonzero chemical potential, the longitudinal propagator is strongly suppressed in the infrared, whereas the transverse one remains comparatively stable Bazavov et al. (2019). Moreover, the transverse channel is usually determined with better numerical precision in lattice simulations, making it the most reliable observable. Thus, the results presented in this work provide the CI model prediction for the transverse component over a broader temperature range than for the longitudinal one, which is calculated only up to $T_{c}$. This extension constitutes a valuable step toward a more complete description, as it allows for analytical expressions of screening masses and BSAs. However, the same analytical treatment also increases the complexity of incorporating longitudinal effects beyond $T_{c}$, making predictions above this value less robust. Nevertheless, the present results offer a consistent baseline that can guide future refinements and extensions of the approach.
@@ III.2 Axial-Vector Mesons
FIG. 6: Longitudinal and transverse screening masses for AV mesons composed of the five quark flavors($\color[rgb]{1,0,0}u$}, $\color[rgb]{0.64,0.0,0.0}d$}, $\color[rgb]{0,0.5,0}s$}, $\color[rgb]{0.55,0.0,0.55}c$} and $\color[rgb]{0,0,1}b$}).The longitudinal components are shown with dashed lines, while the transverse components are shown with solid lines. In the top left panel, the lightest axial-vector meson is shown, the $a_{1}$ meson, The black dashed line corresponds to the free-theory limit, $m=2\pi T/T_{c}$; the top center panel displays the $K_{1}^{*}$ screening mass results; the top right panel displays strangeonium axial-vector mesons. The bottom left panel is dedicated to axial-vector $D$-mesons ($D_{1}$ and $D_{{\color[rgb]{0,0.5,0}s}1}$); the bottom center panel is committed for axial-vector $B$-mesons ($B_{1}$, $B_{{\color[rgb]{0,0.5,0}s}1}$ and $B^{*}_{{\color[rgb]{0.55,0.0,0.55}c}{\color[rgb]{0,0,1}b}}$) and, the bottom right panel shows the results for charmonium and bottomonium axial-vector mesons. Squares and crosses denote lattice results from Refs. Bazavov and Weber (2021); Cheng et al. (2011), inverted triangles represent results from an effective chiral Lagrangian in Ref. Song (1993). Hexagons indicate QCD sum-rule predictions based on finite-temperature meson masses for the heaviest mesons from Ref. Yazici (2016). Although these results do not correspond to screening masses, they provide a useful benchmark in sectors where screening-mass calculations are not yet available.
@@
Figure 6 illustrates the temperature dependence of the screening masses over the range $T/T_{c}\in[0,250]$ MeV. Transverse results are shown with solid lines, longitudinal results with dashed lines. Analogously to the vector mesons, we show the free-theory limit with a black dashed line in the case of light mesons. Our results are compared with lattice data Bazavov and Weber (2021); Cheng et al. (2011); Mukherjee (2009), an effective chiral Lagrangian Song (1993), and QCD sum-rule predictions Yazici (2016). It should be stressed that the results derived from the last two approaches are reliable only for temperatures lower than the critical temperature, for the same reasons explained for V mesons.